# Machine Learning needs its own Randomness Standard: Randomised Smoothing and PRNG-based attacks

### v1

@@ Abstract
Randomness supports many critical functions in the field of machine learning (ML) including optimisation, data selection, privacy, and security. ML systems outsource the task of generating or harvesting randomness to the compiler, the cloud service provider or elsewhere in the toolchain. Yet there is a long history of attackers exploiting poor randomness, or even creating it – as when the NSA put backdoors in random number generators to break cryptography. In this paper we consider whether attackers can compromise an ML system using only the randomness on which they commonly rely. We focus our effort on Randomised Smoothing, a popular approach to train certifiably robust models, and to certify specific input datapoints of an arbitrary model. We choose Randomised Smoothing since it is used for both security and safety – to counteract adversarial examples and quantify uncertainty respectively. Under the hood, it relies on sampling Gaussian noise to explore the volume around a data point to certify that a model is not vulnerable to adversarial examples. We demonstrate an entirely novel attack against it, where an attacker backdoors the supplied randomness to falsely certify either an overestimate or an underestimate of robustness. We demonstrate that such attacks are possible, that they require very small changes to randomness to succeed, and that they can be hard to detect. As an example, we hide an attack in the random number generator and show that the randomness tests suggested by NIST fail to detect it. We advocate updating the NIST guidelines on random number testing to make them more appropriate for safety-critical and security-critical machine-learning applications.
@@ 1 Introduction
Randomness is crucial in machine learning (ML), serving a number of purposes across different areas. One use case is in federated learning, where it helps with user selection to ensure a diverse representation of participants and prevent bias [34]; it can also provide privacy amplification in the process [3]. It is very heavily used in optimisation, providing the foundation for Stochastic Gradient Descent [56], as well as generally for data sampling, enabling the selection of representative subsets for model training [47]. It enables Monte Carlo methods [48]. It is essential in active learning, a process where an algorithm actively selects the most informative samples for labelling, in order to reduce the labelling cost [31]. It forms the basis of differential privacy, the de facto standard for quantifying privacy in ML, where noise is added to data in such a way as to protect individual privacy while still allowing for accurate analysis [20]. It contributes to the generation of synthetic data, expanding the training set and enhancing the generalisation capabilities of machine learning models [36]. In short, randomness is widely used and highly significant for ML. Yet little thought has been given to the vulnerabilities that might result from weak randomness.
@@
We therefore explore the extent to which an attacker can change the safety-critical decision-making of a target system, purely by exploiting or tinkering with random number generators. We focus our efforts on Randomised Smoothing, a standard technique for quantifying uncertainty [^1] The literature does not view Randomised Smoothing as an uncertainty estimation technique, yet it is one when it is used to certify a prediction around a target datapoint. in a given blackbox model prediction [13].This is heavily used in practice to combat adversarial examples; it can even be used to provide robustness certification. It samples isotropic Gaussian noise and adds it to a critical datapoint, in order to measure the probability that the addition of noise may cause the model prediction to change, leading to a spurious decision. Sampling high-quality noise is crucial for this purpose, yet in practice we rarely check how normal our Gaussian noise is.
@@
In this paper, we construct two attacks against Randomised Smoothing that assume an attacker can influence the random number generator on which the model relies. The first is a naive attack that simply replaces a Gaussian distribution with a different distribution e.g. Laplace noise. This disrupts confidence quantification significantly for both over- or under-estimation, but is relatively easy to detect. We present the attack intuition visually in Figure 1. Following this naive proof of concept, we construct a more powerful and covert attack: a bit-flipping PRNG attacker that changes only a single bit out of every 64 bits deep in the implementation of the random number generator. We show that this can cause mis-quantification of the true confidence by up to a factor of ×$\times 81$ and is extremely hard to detect. This change to the PRNG is covert in that it does not cause it to fail the official NIST randomness suite of tests.
@@
We argue that similar randomness-based attacks can be devised against other ML techniques, such as differential privacy. It follows that the standards are insufficient to guarantee the security and privacy of machine learning systems in the face of sophisticated adversaries. We then discuss practical ways of tackling these vulnerabilities in practice. We aim to empower researchers and practitioners to understand the extent to which they place their trust in their toolchain’s source of randomness, and develop more robust defences against the abuse of this trust.
@@ 2 Related work
Figure 1: A visual example of how manipulated randomness can result in mis-perception of real model confidence. A normal Randomised Smoothing procedure is shown on the left – the original datapoint $d$ is perturbed to get points in the dotted region, ultimately deriving a true $\epsilon$ confidence region. The attacker biases the sampled noise and changes the estimation of Randomised Smoothing, as shown on the right. This overestimates the confidence and leads to an incorrect prediction. Noise can also be biased to reduce the $\epsilon$ instead.
In this section we cover the work that relates to the attack we developed. Section 2.1 covers adversarial examples and Section 2.3 discusses how Randomised Smoothing can use randomness to counteract them. It is followed by a section on randomness.
@@ 2.1 Adversarial Examples
Research into adversarial examples took off in 2014 when Szegedy et al. [63] discovered that deep learning models are very susceptible to such attacks. This led to the formulation of an evasion attack as an optimisation problem by Biggio et al. [6]. To determine the optimal adversarial perturbation on a given input $x^{0}$, an attacker aims to minimise $L:\ \mathcal{X}\rightarrow\mathbb{R}$, where $L$ is the model’s discriminant function:
@@ 2.2 Certification in Machine Learning
Certification techniques attempt to invert Equation 1 to determine the maximum perturbation radius around an input, inside which no adversarial examples exist. This is different from what the security community means by certification in the context of formal verification [16]: this refers to programmatic proof that a piece of software performs exactly as intended, matching some high-level abstraction. For example, a certified compiler will always correctly translate legal code correctly into its compiled binary or throw an appropriate error [39]. If the high-level abstraction of a machine learning model is considered to be the function $f:\mathcal{X}\rightarrow\mathcal{Y}$, mapping inputs to their true classes, then certification here would refer to making sure that, for a given volume around a point from $\mathcal{X}$, the prediction does not change from $\mathcal{Y}$. Many certification techniques in current use are approximate and, despite the fact that theorems may be offered around their behaviour, they only provide probabilistic guarantees for the absence of adversarial examples in an $\epsilon$ ball around a given input [14].
Randomised Smoothing is not the only certification technique. Other mechanisms exist, e.g. by constructing a bounding polytope instead of a hypersphere, or using either linear relaxation or interval bound propagation [69]. However, these require changes to the training process and model re-engineering, which in turn makes them hard to scale up to more complex model architectures and higher dimensional data [16]. Therefore, Randomised Smoothing is considered the state-of-the-art certification algorithm for machine learning models. It can be applied to any underlying model in a black-box manner, and scaled up to larger model architectures. It is used both for security, i.e. to combat adversarial examples, and safety, i.e. to provide confidence in predictions. That is why we made it the target of this work.
@@ 2.3 Randomised Smoothing
This formulation allows easy approximation of $g$ for a given confidence bound using a Monte Carlo algorithm by sampling $\epsilon$ from $\mathcal{N}(0,\sigma^{2}\mathit{I})$ [13]. While previous work used differential privacy [19, 37] and Renyi divergence [66, 40] to determine the radius of the hypersphere around an input inside which the absence of adversarial examples can be provably verified, Cohen et al. [13] used the Neyman-Pearson lemma [50] to obtain a certification radius which is also provably maximal. The certified radius obtained from randomised smoothing is:
@@ 2.4 Randomness in (Adversarial) ML
Randomness is heavily used in machine learning and there are many applications. Stochastic Gradient Descent, the randomised optimisation algorithm that arguably enabled modern deep learning, relies on randomness for data sampling [56]; randomised dropout improves generalisation [22]; active learning approaches rely on biased samplers to enable faster learning [31]; randomised transformations are used to introduce trained invariance [41]; while randomised sampling of data can lead to improved privacy[3], and even bound privacy attacks [65].
Adversarial ML uses randomness in both attack and defense. Adversarial examples benefit significantly from random starts [14]; while many ML system enginners advocate for defences based on random pre-processing, where inputs are stochastically transformed before inference [52, 23, 2]. We are aware of only two attacks so far that only use randomness: the batch reordering attack [62] and the randomised augmentation attack [54], where the order of data and the randomness of the augmentation are manipulated respectively to introduce backdoors into a target model.
@@ 3.1 Random Number Generators
While many distributions can be sampled to generate random sequences, in the context of computer programming, generally the uniform distribution $\mathcal{U}(a,b)$ is used, i.e. every number between $a$ and $b$ has an equal probability of selection. As discussed later in Section 3.3, the standard uniform distribution ($\mathcal{U}(0,1)$) can be transformed into any other random distribution, and is therefore a natural starting point for random number generators (RNG). There are quite a few sources of entropy that a computer can rely on to generate random numbers such as the timing of keystrokes or mouse movements [21]. For example, consider an RNG relying on timing of keystrokes as the source of entropy. Generating one random bit from this RNG can be modelled as an experiment phrased as “Is the time taken between two keystrokes in milliseconds an even number?". The outcome of this experiment can either be 0 or 1, thereby generating a random bit. The function that assigns a real value to each outcome of a random experiment such as the one described here is called a random variable [57]. The distribution function of a 1D random variable $X$ over the real line can be defined as
@@
The probability mass function $p$ for a discrete random variable $X$ is
$p(x)=\mathbb{P}\{X=x\}.$ (5)
@@
These issues led to the development of pseudo-random number generators (PRNGs) which rely on deterministic calculations to generate sequences of numbers that are not truly random, yet appear to be. Von Neuman proposed the first PRNG [68] in 1946 using the middle-square method. PRNGs use the current number in the sequence or some other counter to generate the next number, but they need a starting point – a seed – to generate the first number or counter value. The sequence of numbers is thus a function of the seed, and two instances of the same PRNG will produce the same sequence if the seeds are identical [^2] Not all PRNGs work like this, but this is a desirable property to ensure reproducibility. Cryptographic PRNGs typically combine hardware and pseudorandom components in such a way that both have to fail to make key material easily predictable by an opponent.. The following sections present the process of generating random numbers, and transforming them to sample from the normal distribution. This is followed by an overview of statistical tests to determine the quality of random numbers produced by a PRNG, and finally a discussion of possible attacks.
@@ 3.2 The Linear Congruential Method and PCG64
Equation 6 can be used to generate random numbers between 0 and $m$. Different choices of $a$ and $c$ affect the performance. The permuted congruential generator (PCG) [51] is widely used in ML and uses an adaptation of this method. The 64-bit version has a 128-bit state variable, which is advanced according to Equation 6, i.e. $X_{n}$ gives the $n$th state and is a 128-bit integer. A 64-bit random number can be generated from this 128-bit state as:
output = rotate64(((uint64_t) state ^ (uint64_t)(state>>64)), state>>122)
@@ 3.3 Sampling from a Normal Distribution
The numbers generated by PCG64 can be assumed to be sampled from the discrete uniform distribution $\mathcal{U}(0,2^{64})$. However, most randomised algorithms need this to be transformed into a different distribution, usually $\mathcal{U}(0,1)$ or $\mathcal{N}(0,1)$, which can then be transformed into any uniform or normal distribution trivially. Transforming discrete $\mathcal{U}(0,2^{64})$ to continuous $\mathcal{U}(0,1)$ can be done by simply discarding the first 11 bits of the random number [^4] These bits contain information about the exponent in 64-bit floating point numbers [1]., typecasting as double, and dividing by $2^{53}$ to effectively shift the decimal point from after the least significant bit to before the most significant bit. Analytically, a $\mathcal{U}(0,1)$ sample can be transformed to correspond to a sample from any other distribution with a distribution function $F$ defined in Equation 4 by inverting it:
$x^{\prime}=F^{-1}(x),\ x\in(0,1).$ (7)
However, for some distributions, including the normal distribution, it is impossible to invert $F$ analytically, and other means must be employed. The rectangle-wedge-tail method, first proposed by Marsaglia in 1961 [44, 46, 45], is an ingenious way to compute the absolute value of a normally distributed random number.
A simplified version is presented here, and the reader is directed to Knuth’s ‘The Art of Computer Programming’ [33] for a more detailed discussion. The normal distribution is split into 32 regions, and the first 8 bits of the random number are used to determine which region the target number should fall in. Every region can be thought of as the uniform distribution lying between 0 and an upper bound, which increases with every subsequent region (see Figure 5) These uniform distributions, when stacked on top of each other, approximate the normal distribution. Within each region, 55 bits of the random number are transformed into a float lying between zero and the upper bound by multiplying with a constant stored separately. The last bit is used to assign a sign to the target number. This method can efficiently generate a normally distributed number 99.3% of the time. If the target number is determined to fall in the outermost region, the polar method [9] can be used. How the random integer generated by PCG64 is transformed to a normally distributed random floating point number will come into play when the attack is discussed below.
{tcolorbox}
[colback=black!5!white,colframe=black,title=Attack Scenario Example]
Example Setting: Detection of an enemy tank is being performed from a satellite image. To make sure that the prediction is not a spurious correlation and to counteract camouflage paint, Randomised Smoothing is used. PRNG generates the noise for Randomised Smoothing. Randomness can come from the inference platform i.e. the cloud hardware or ML-as-a-Service API; alternatively the user can apply noise themselves [^5] Note that this may result in quantisation artefacts, e.g. a pixel with value 245 noised to 235.3 gets quantised to 236, and potentially changes the performance of Randomised Smoothing. .
A Normal User: A user attempting to certify the robust l2 radius that does not contain adversarial examples around an input in the above setting. Here, our user may be an analyst who aims to find enemy tanks to target with missiles.
Attacker: An adversary with the objective of manipulating the certified radius obtained by the user. An attacker may increase it to make the prediction appear more robust than it actually is, i.e. convince an the user that tank is present at a given location, or decrease it to make the user uncertain about their prediction or even abstain from it, i.e. force them to not attack a tank at a given location.
Defender: The defender’s goal is to detect the presence of a randomness-based adversary to determine if the noise function used for certification deviates significantly from white noise or not.

### v2

# Machine Learning needs Better Randomness Standards: Randomised Smoothing and PRNG-based attacks

@@ Abstract
Randomness supports many critical functions in the field of machine learning (ML) including optimisation, data selection, privacy, and security. ML systems outsource the task of generating or harvesting randomness to the compiler, the cloud service provider or elsewhere in the toolchain. Yet there is a long history of attackers exploiting poor randomness, or even creating it – as when the NSA put backdoors in random number generators to break cryptography. In this paper we consider whether attackers can compromise an ML system using only the randomness on which they commonly rely. We focus our effort on Randomised Smoothing, a popular approach to train certifiably robust models, and to certify specific input datapoints of an arbitrary model. We choose Randomised Smoothing since it is used for both security and safety – to counteract adversarial examples and quantify uncertainty respectively. Under the hood, it relies on sampling Gaussian noise to explore the volume around a data point to certify that a model is not vulnerable to adversarial examples. We demonstrate an entirely novel attack, where an attacker backdoors the supplied randomness to falsely certify either an overestimate or an underestimate of robustness for up to 81 times. We demonstrate that such attacks are possible, that they require very small changes to randomness to succeed, and that they are hard to detect. As an example, we hide an attack in the random number generator and show that the randomness tests suggested by NIST fail to detect it. We advocate updating the NIST guidelines on random number testing to make them more appropriate for safety-critical and security-critical machine-learning applications.
@@ 1 Introduction
Randomness is crucial in machine learning (ML), serving a number of purposes across different areas. One use case is in federated learning, where it helps user selection to ensure a diverse representation of participants and prevent bias [34]; it can also provide privacy amplification in the process [3]. It is heavily used in optimisation, providing the foundation for Stochastic Gradient Descent [54], as well as generally for data sampling, enabling the selection of representative subsets for model training [45]. It enables Monte Carlo methods [46]. It is essential in active learning, a process where an algorithm actively selects the most informative samples for labelling, in order to reduce the training cost [31]. It forms the basis of differential privacy, the de facto standard for quantifying privacy, where noise is added to data in such a way as to protect individual privacy while still allowing for accurate analysis [19]. It contributes to the generation of synthetic data, which aids in expanding the training set and enhancing the generalisation capabilities of ML models [36]. In short, randomness is widely used and highly significant for ML.
@@
Yet little thought has been given to the vulnerabilities that might result from weak randomness. We therefore explore the extent to which an attacker can change the safety-critical decision-making of a target system, purely by exploiting or tinkering with random number generators. We focus our efforts on Randomised Smoothing, a standard technique for quantifying uncertainty [^1] The literature does not view Randomised Smoothing as an uncertainty estimation technique, yet it is one when it is used to certify a prediction around a target datapoint. in a given blackbox model prediction [12]. This is heavily used in practice to combat adversarial examples; it can even be used to provide robustness certification. It samples isotropic Gaussian noise and adds it to a critical datapoint in order to measure the probability that the addition of noise causes the model prediction to change, leading to a spurious decision . Sampling high-quality noise is crucial for this purpose, yet in practice we rarely check how normal our Gaussian noise is.
@@
In this paper, we construct two attacks against Randomised Smoothing that assume an attacker can influence the random number generator on which the model relies. The first is a naive attack that simply replaces a Gaussian distribution with a different distribution e.g. Laplace noise. This disrupts confidence quantification significantly for both over- or under-estimation, but is relatively easy to detect. We present the attack intuition visually in Figure 1. Following this naive proof of concept, we construct a more powerful and covert attack: a bit-flipping PRNG attacker that changes only a single bit out of every 64 bits deep in the implementation of the random number generator. We show that this can cause mis-quantification of the true confidence by up to a factor of ×$\times 81$ and is extremely hard to detect. This change to the PRNG is covert, in that it does not cause it to fail the official NIST suite of randomness tests.
@@
This highlights the inadequacy of current standards and defences in protecting against attacks targeting randomness in machine learning. We argue that similar randomness-based attacks can be devised against other ML techniques, such as differential privacy. It follows that the standards are insufficient to guarantee the security and privacy of machine learning systems in the face of sophisticated adversaries. We then discuss practical ways of tackling these vulnerabilities. We aim to empower researchers and practitioners to understand the extent to which they place their trust in their toolchain’s source of randomness, and develop more robust defences against the abuse of this trust.
@@ 2 Related work
Figure 1: A visual example of how manipulated randomness affects Randomised Smoothing. Here we aim to certify a point in the middle. To do that we sample data around it using noise – a normal noise sampling for Randomised Smoothing is shown in black. Manipulated random noise with lower kurtosis, skewness and higher kurtosis are shown in order from left to right. Manipulated noise misrepresents the true decision space and leads to incorrect prediction confidence.
In this section we cover the work that relates to the class of attacks we developed. Section 2.1 covers adversarial examples and Section 2.3 discusses how Randomised Smoothing can use randomness to counteract them for both safety and security. It is followed by a section on randomness.
@@ 2.1 Adversarial Examples
The evasion attack problem was formulated as an optimisation problem by Biggio et al. [6] in 2013. Szegedy et al. [61] demonstrated that deep learning models are indeed highly susceptible to evasion attacks. To determine the optimal adversarial perturbation on a given input $x^{0}$, an attacker aims to minimise $L:\ \mathcal{X}\rightarrow\mathbb{R}$, where $L$ is the model’s discriminant function [6]:
@@ 2.2 Certification in Machine Learning
Certification techniques attempt to invert Equation 1 to determine the maximum perturbation radius around an input, inside which no adversarial examples exist. This is different from what the security community means by certification in the context of formal verification [15]: that refers to programmatic proof that a piece of software performs exactly as intended, matching some high-level abstraction. For example, a certified compiler will always correctly translate legal code correctly into its compiled binary or throw an appropriate error [39]. If the high-level abstraction of a machine learning model is considered to be the function $f:\mathcal{X}\rightarrow\mathcal{Y}$, mapping inputs to their true classes, then certification here would imply that, for a given volume around a point from $\mathcal{X}$, the prediction does not change from $\mathcal{Y}$. Many certification techniques in current use are approximate and, despite the fact that theorems may be offered around their behaviour, they only provide probabilistic guarantees for the absence of adversarial examples in an $\epsilon$ ball around a given input [13].
Randomised Smoothing is not the only certification technique. Other mechanisms exist, e.g. by constructing a bounding polytope instead of a hypersphere, or using either linear relaxation or interval bound propagation [67]. However, these require changes to the training process and model re-engineering, which in turn makes them hard to scale up to more complex model architectures and higher dimensional data [15]. Randomised Smoothing is one of the most widely used certification techniques and it underpins current state-of-the-art certification algorithm for machine learning models [68]. It can be applied to any underlying model in a black-box manner and scaled up to larger model architectures. It is used both for security, i.e. to combat adversarial examples, and safety, i.e. to provide confidence in predictions. That is why we made it the target of this work.
@@ 2.3 Randomised Smoothing
This formulation allows for easy approximation of $g$ for a given confidence bound using a Monte Carlo algorithm by sampling $\epsilon$ from $\mathcal{N}(0,\sigma^{2}\mathit{I})$ [12]. While previous work used differential privacy [18, 37] and Renyi divergence [64, 40] to determine the radius of the hypersphere around an input inside which the absence of adversarial examples can be provably verified, Cohen et al. [12] used the Neyman-Pearson lemma [48] to obtain a certification radius which is also provably maximal. The certified radius obtained from randomised smoothing is:
@@ 2.4 Randomness in (Adversarial) ML
Randomness is heavily used in machine learning. Stochastic Gradient Descent, the randomised optimisation algorithm that arguably enabled modern deep learning, relies on randomness for data sampling [54]; randomised dropout improves generalisation [21]; active learning approaches rely on biased samplers to enable faster learning [31]; randomised transformations are used to introduce trained invariance [41]; while randomised sampling of data can lead to improved privacy[3], and even bound privacy attacks [63].
Adversarial ML uses randomness in both attack and defense. Adversarial examples benefit significantly from random starts [13], while many ML system enginners advocate defences based on random pre-processing where inputs are stochastically transformed before inference [50, 22, 2]. We are aware of only two attacks so far that depend on exploiting randomness: the batch reordering attack [60] and the randomised augmentation attack [52], where the order of data and the randomness of the augmentation are manipulated respectively to introduce backdoors into a target model.
@@ 3.1 Random Number Generators
While many distributions can be sampled to generate random sequences, in the context of computer programming, generally the uniform distribution $\mathcal{U}(a,b)$ is used, i.e. every number between $a$ and $b$ has an equal probability of selection. As discussed later in Section 3.3, the standard uniform distribution ($\mathcal{U}(0,1)$) can be transformed into any other random distribution and is therefore a natural starting point for random number generators (RNG). There are quite a few sources of entropy that a computer can use to generate random numbers such as the timing of keystrokes or mouse movements [20]. For example, consider an RNG relying on timing of keystrokes as the source of entropy. Generating one random bit from this RNG can be modelled as an experiment phrased as “Is the time taken between two keystrokes in milliseconds an even number?". The outcome of this experiment can either be 0 or 1, thereby generating a random bit. The function that assigns a real value to each outcome of a random experiment such as the one described here is called a random variable [55]. The distribution function of a 1D random variable $X$ over the real line can be defined as
@@
These issues led to the development of pseudo-random number generators (PRNGs) which rely on deterministic calculations to generate sequences of numbers that are not truly random, yet appear to be. Von Neuman [66] proposed the first PRNG in 1946 using the middle-square method. PRNGs use the current number in the sequence or some other counter to generate the next number, but they need a starting point – a seed – to generate the first number or counter value. The sequence of numbers is thus a function of the seed, and two instances of the same PRNG will produce the same sequence if the seeds are identical [^2] Not all PRNGs work like this, but this is a desirable property to ensure reproducibility. Cryptographic PRNGs typically combine hardware and pseudorandom components in such a way that both have to fail to make key material easily predictable by an opponent.. The following sections present the process of generating random numbers and transforming them into samples from the normal distribution. This is followed by an overview of statistical tests to determine the quality of random numbers produced by a PRNG, and finally a discussion of possible attacks.
@@ 3.2 The Linear Congruential Method and PCG64
Equation 5 can be used to generate random numbers between 0 and $m$. Different choices of $a$ and $c$ affect the performance. The permuted congruential generator (PCG) [49] is widely used and uses an adaptation of this method. The 64-bit version has a 128-bit state variable, which is advanced according to Equation 5, i.e. $X_{n}$ gives the $n$th state and is a 128-bit integer. A 64-bit random number can be generated from this 128-bit state as:
output=rotate64(state^(state>>64)),state>>122)
@@ 3.3 Sampling from a Normal Distribution
The numbers generated by PCG64 can be assumed to be sampled from the discrete uniform distribution $\mathcal{U}(0,2^{64})$. However, most randomised algorithms need this to be transformed into a different distribution, usually $\mathcal{U}(0,1)$ or $\mathcal{N}(0,1)$, which can then be transformed into any uniform or normal distribution trivially. Transforming discrete $\mathcal{U}(0,2^{64})$ to continuous $\mathcal{U}(0,1)$ can be done by simply discarding the first 11 bits of the random number [^4] These bits contain information about the exponent in 64-bit floating point numbers [1], typecasting as double, and dividing by $2^{53}$ to effectively shift the decimal point from after the least significant bit to before the most significant bit.
The Ziggurat method [44] is state-of-the-art when it comes to sampling from decreasing densities such as the normal distribution. At a high level, this involves dividing the distribution into smaller pieces, choosing one of these pieces randomly, and then sampling from it. The smaller pieces into which the distribution is divided are chosen such that the areas under the density curve are equal, i.e. each piece is equally likely (shown in Figure 2). The lower 8 bits (index bits) are used to select one of these pieces at random, the next bit (sign bit) is used to randomly assign a sign, and finally, the remaining bits (distribution bits) are used to sample a float from the uniform distribution corresponding to the rectangular area selected using the index bits.
Figure 2: Sampling from a Gaussian distribution constructed using Uniform distributions with the Ziggurat Method [44].